solve each system of equations using the substitution method 11x+16y=922x+3y=18
Answer by mananth(16949)

11x+16y= 92.............1
3y=18-2x
/3
y=6-0.67 x
Plug the value of y in (1)
11x+16(6 -0.67x)= 92
11x+96-10.67x=92
11x-10.67x=92-96
0.33x=-4
/0.33
x=-12
Plug the value of x in (1)
11x+16y= 92
11*-12+16y=92
-132+16y =92
16y=224
/16
y=14

CHECK
3 y = 18 + -2 x
42 = 18 + 24